OpenStudy (anonymous):

simplify completly 14x^5y^4 + 21x^3y^2 over 7x^3y

OpenStudy (callisto):

( 14x^5y^4 + 21x^3y^2 ) / ( 7x^3y) = 7x^3y^2(2x^2+3) /( 7x^3y) =y(2x^2+3)

OpenStudy (anonymous):

not right answer :(

OpenStudy (callisto):

sorry.. made a mistake in the second step! = 7x^3y^2(2x^2y^2+3) /( 7x^3y) = y(2x^2y^2+3) = 2x^2y^3 +3y